In a stunt being filmed for a movie, a sports car overtakes a truck towing a ramp, drives up and off the ramp, soars into the air, and then lands on top of a flat trailer being towed by a second truck. The tops of the ramp and the flat trailer are the same height above the road, and the ramp is inclined 11° above the horizontal. Both trucks are driving at a constant speed of 18 m/s, and the flat trailer is 13 m from the end of the ramp. Neglect air resistance, and assume that the ramp changes the direction, but not the magnitude, of the car's initial velocity. What is the minimum speed the car must have, relative to the road, as it starts up the ramp?
Answer #1

here,

theta = 11 degree

final speed , v = 18 m/s

s = 13 m

accelration , a = - g * sin(theta)

a = - 9.8 * sin(11) = - 1.87 m/s^2

let the initial speed be u

v^2 - u^2 = 2 * a * s

18^2 - u^2 = - 2 * 1.87 * 13

u = 19.3 m/s

the initial speed of car must be 19.3 m/s